# Register Any Point: Scaling 3D Point Cloud Registration by Flow Matching

### v1

@@ 3.4 Canonicalized Registration Pipeline
Directly generating millions of points is computationally inefficient and unstable. Instead, we construct a compact, geometry-aware representation by sampling keypoints in each view and attaching local descriptors. For each input point cloud $\mathbf{P}_{i}$, we first perform voxel downsampling with voxel size $v_{d}$, obtaining a reduced cloud $\mathbf{P}_{i}^{v}$. We then apply the farthest point sampling (FPS) to $\mathbf{P}_{i}^{v}$ to select $K_{i}$ keypoints, forming $\mathbf{Q}_{i}\in\mathbb{R}^{3\times K_{i}}$, with the total keypoint count $K:=\sum_{i}K_{i}$ across all views. To obtain uniform coverage over the scene, we choose $K_{i}$ in proportion to the metric scale of $\mathbf{P}_{i}$.
@@
To encode local geometry, for each sampled point in $\mathbf{Q}_{i}$ we define a local patch by a ball query of radius $r_{s}=20v_{d}$ in the reduced point cloud $\mathbf{P}_{i}^{v}$ and extract a local descriptor from the normalized points within this patch. We use the lightweight, rotation-invariant MiniSpinNet [ao2023cvpr-buffer, seo2025iccv] pretrained on 3DMatch [zeng2017cvpr], yielding $\mathbf{F}_{i}\in\mathbb{R}^{32\times K_{i}}$. Concatenating all views, we obtain sampled points $\mathcal{Q}=\{\mathbf{Q}_{i}\in\mathbb{R}^{3\times K_{i}}\mid i=1,\ldots,N\}$ and their local features $\mathcal{F}=\{\mathbf{F}_{i}\in\mathbb{R}^{D\times K_{i}}\mid i=1,\ldots,N\}$, which serve as a compressed yet informative representation of the dense input clouds $\mathcal{P}$.
Canonicalization of inputs and targets.
@@
To make training invariant to global pose and metric scale, we canonicalize both the conditioning representation and the flow target in a shared frame. For each view $i$, we first translate the sampled points $\mathbf{Q}_{i}$ so that its center of mass is at the origin. We then compute a global scale factor $s$ as the longest edge length of the bounding box of the view with the most points, and scale all centered point sets by $1/s$, so that the entire scene fits in a unit cube. Finally, we apply a random 3D rotation to each centered, scaled cloud. This yields normalized unposed keypoints $\bar{\mathcal{Q}}=\{\bar{\mathbf{Q}}_{i}\in\mathbb{R}^{3\times K_{i}}\mid i=1,\dots,N\}$ and corresponding similarity transforms $\bar{\mathbf{T}}_{i}\in\mathrm{SIM}(3)$ that map the original $\mathbf{Q}_{i}$ to $\bar{\mathbf{Q}}_{i}$.
@@
The training target is defined in the same canonical frame. We first transform the keypoints by the ground-truth poses $\mathcal{T}$ to obtain registered keypoints $\mathcal{Q}^{r}=\{\mathbf{Q}_{i}^{r}\}_{i=1}^{N}$ and merge them into $\mathbf{Q}^{r}=\bigcup_{i=1}^{N}\mathbf{Q}_{i}^{r}$. We then (i) recenter $\mathbf{Q}^{r}$ at the origin, (ii) apply the same random rotation used for the view with the most points to fix a reference orientation, and (iii) scale by the global factor $s$, resulting in the normalized registered point cloud $\bar{\mathbf{Q}}^{r}$, which is set as the target $\mathbf{X}(0)$.
Conditional flow model.
@@
We adopt the Diffusion Transformer [peebles2023iccv] for $\mathbf{V}_{\theta}$, and, following VGGT [wang2025cvpr-vggt], employ a transformer with alternating-attention blocks (Fig. 2). Specifically, we alternate per-view self-attention within each point cloud, which consolidates view-specific structure, with global attention over all point tokens to fuse information across views. Our model comprises $L=10$ alternating-attention blocks with hidden dimension $d=512$ and $h=8$ attention heads, totaling 73 million parameters.
@@
The model’s condition $\mathbf{C}=f_{\text{emb}}(\bar{\mathcal{Q}},\mathcal{F},s)$ is obtained via a linear feature embedder $f_{\text{emb}}$ with the concatenated local descriptors $\mathcal{F}$, positional encodings of the normalized coordinates $\bar{\mathcal{Q}}$, and a global embedding of the scene scale $s$. The flow network $\mathbf{V}_{\theta}$ takes $\mathbf{X}(t)$ and $\mathbf{C}$ as input and is trained with the conditional flow matching loss [lipman2023iclr].
Lifting to dense registered point clouds.
@@
At inference, the model generates a canonical registered keypoint cloud $\hat{\mathbf{X}}(0)$. Using the rigidity-forcing procedure in Sec. 3.3, we recover per-view rigid transformations $\hat{\mathbf{T}}_{i}$ that align $\bar{\mathbf{Q}}_{i}$ to the corresponding subsets of $\hat{\mathbf{X}}(0)$. The overall transformation from the original input frame of view $i$ to the final registered frame is $\mathbf{T}_{i}=\hat{\mathbf{T}}_{i}\bar{\mathbf{T}}_{i}$, which we apply to all points in the dense cloud $\mathbf{P}_{i}$. Finally, we undo the global scaling by $s$ to obtain the registered point clouds at the metric scale.
@@ 4 Experimental Evaluation
The main focus of this work is a scalable generative model that aligns multiple point clouds using flow matching while enforcing rigidity constraints. We present our experiments on pairwise and multi-view point cloud registration to show the capabilities of our method.
@@ 4.1 Experimental Setup
Implementation details. We train our model in two phases. In phase one, we train for four days using training samples with point cloud view count $N$ from two to six. In phase two, we do not limit the maximum view count and keep training for three more days. We use 8 NVIDIA A100 GPUs with 80GB VRAM each. We use Muon [liu2025arxiv-muon] as the optimizer with an initial learning rate of $2\cdot 10^{-3}$ for matrix-like parameters and $2\cdot 10^{-4}$ for vector-like parameters.
Unlike VGGT [wang2025cvpr-vggt] or RPF [sun2025neurips], our training samples have varying numbers of views $N$ and point-token counts $K$. To efficiently train under this irregular data setup, we devise a dynamic batching strategy that allocates samples to each GPU with a suitable batch size. We set the maximum token count per batch on one GPU to 100,000 in phase one and 80,000 in phase two to fill the GPU memory.
Table 1: Datasets used for training our model.
Dataset #Scenes #Scans Real/Synth. #Train
KITTI [geiger2012cvpr] 22 16,453 Real 19
KITTI360 [liao2022pami] 9 21,255 Real 7
Apollo [huang2018cvprws] 11 25,036 Real 9
nuScenes [caesar2020cvpr] 642 24,320 Real 642
MulRAN [kim2020icra] 4 4,477 Real 3
Boreas [burnett2023ijrr-boreas] 2 3,613 Real 2
Oxford Spires [tao2025ijrr-oxfordspires] 6 5,781 Real 5
VBR [brizi2024icra-vbr] 5 8,647 Real 4
UrbanNav [hsu2023navi-urbannav] 4 4,228 Real 3
WildPlace [knights2023icra-wildplaces] 2 2,613 Real 2
HeLiPR [jung2024ijrr] 3 10,882 Real 3
KITTI-Carla [deschaud2021arxiv-kitticarla] 7 7,729 Synth. 5
3DMatch [zeng2017cvpr] 82 29,314 Real 48
NSS [sun2025jprs] 6 54,396 Real 3
ScanNet [dai2017cvpr] 661 75,299 Real 511
ScanNet++ [yeshwanth2023iccv-scannetpp] 220 88,783 Real 194
ModelNet [wu2015cvpr] 12,308 24,616 Synth. 9,800
@@
For our model, we curate over 100k samples from 17 diverse datasets spanning outdoor, indoor, and object-centric settings as summarized in Tab. 1. Each sample consists of $N$ point clouds, where $2\leq N\leq 24$. We split the samples into training and validation sets with an approximate ratio of 9:1 while we also exclude sequences used for testing in common registration benchmarks from the training set.
The datasets span diverse scenes across multiple continents, captured by LiDAR and depth cameras with varying resolutions and fields of view. We curate both single-frame and sequence-accumulated submap samples, ensuring sufficient overlap between point clouds in each sample.
Testing data. For pairwise registration evaluation, we build our benchmark based on the established generalizability benchmark [seo2025iccv]. The benchmark includes 10 datasets: 3DMatch [zeng2017cvpr], 3DLoMatch [huang2021cvpr-predator], NSS [sun2025jprs], TIERS [sier2023rs-tiers], KITTI [geiger2012cvpr], Waymo [sun2020cvpr], KAIST [jung2024ijrr], MIT [tian2023iros-kimeramultidata], ETH [pomerleau2012ijrr], NCD [ramezani2020iros] with different scales and sensor modalities, where TIERS, Waymo, KAIST, MIT, and NCD are zero-shot testing datasets for our model.
For multi-view registration evaluation, we follow prior work [gojcic2020cvpr, wang2023cvpr-sghr] and use the common multi-view registration benchmarks on 3DMatch [zeng2017cvpr] and ScanNet [dai2017cvpr]. As these datasets are both indoor depth camera datasets, we also add two zero-shot outdoor LiDAR datasets: Waymo [sun2020cvpr] for scan-level registration and KAIST [jung2024ijrr] for map-level multi-view registration evaluation.
Table 2: Quantitative comparison of the pairwise point cloud registration performance in terms of success rate (%) on 10 diverse datasets. Best results are shown in bold, second best are underscored.
Environment Indoor Outdoor
Dataset 3DMatch 3DLoMatch NSS TIERS KITTI Waymo KAIST MIT ETH NCD
Conven- tional FPFH [rusu2009icra] + FGR [zhou2016eccv] 62.53 15.42 30.82 80.60 98.74 100.00 89.80 74.78 91.87 99.00
FPFH [rusu2009icra] + Quatro [lim2022icra] 8.22 1.74 - 86.57 99.10 100.00 91.46 79.57 51.05 91.03
FPFH [rusu2009icra] + TEASER [yang2020tro] 52.00 13.25 - 73.13 98.92 100.00 89.20 71.30 93.69 99.34
Deep learning-based FCGF [choy2019iccv] 88.18 40.09 42.86 80.11 98.92 97.69 93.55 93.04 55.53 95.68
Predator [huang2021cvpr-predator] 90.60 62.40 92.99 75.74 99.82 86.92 87.09 79.56 54.42 93.68
GeoTransformer [qin2022cvpr-geotransformer] 92.00 75.00 55.59 92.99 99.82 89.23 91.86 95.65 71.53 97.01
BUFFER [ao2023cvpr-buffer] 92.90 71.80 - 88.96 99.64 100.00 97.24 95.65 99.30 99.00
PARENet [yao2024iccv-parenet] 95.00 80.50 - 75.06 99.82 92.31 86.44 84.78 69.42 93.36
BUFFER-X [seo2025iccv] 95.58 74.18 - 93.45 99.82 100.00 99.15 97.39 99.72 99.67
Ours ($S=5$ selected) 95.64 78.22 96.18 95.09 100.00 99.47 99.15 99.13 98.04 99.34
Ours ($S=1$ average) 95.27 77.43 96.39 94.75 100.00 98.51 99.05 98.96 98.30 99.34
Evaluation metrics. We follow previous works on point cloud registration to evaluate the performance of our model using the registration success rate (%) calculated with thresholds on correspondence RMSE for 3DMatch and on translation and rotation error for the other datasets. We use threshold settings of previous works [zeng2017cvpr, sun2025jprs, seo2025iccv], detailed in the supplementary material. For multi-view registration, we follow prior work [gojcic2020cvpr, wang2023cvpr-sghr] reporting the mean rotation error (RE) and translation error (TE). We additionally use the Chamfer distance (CD) to evaluate the quality of the registered point cloud by comparing it with the ground truth aggregated point cloud.
4.2 Results
Pairwise registration evaluation. Tab. 2 shows that our model (with and without rigidity-based generation selection) achieves on par or better performance compared to state-of-the-art methods [seo2025iccv, qin2022cvpr-geotransformer, yao2024iccv-parenet] for pairwise registration on standard benchmarks.
To further demonstrate our model’s robustness to low overlap between point clouds, we follow EYOC [liu2024cvpr-eyoc] to curate testing data on KITTI and Waymo with increasing spatial distance between point clouds (thus decreasing overlap). As shown in Fig. 3, our model shows superior performance over state-of-the-art methods with increasing scan distance.
Multi-view registration evaluation. We evaluate on sparse-view subsets (with $3\leq N\leq 12$) of 3DMatch and ScanNet, with a minimum overlap ratio of $0.2$. We evaluate the generalization ability of our model in outdoor scenarios with Waymo for scan-level registration and KAIST for map-level multi-view registration, also with the number of input views set to $3\leq N\leq 12$. Comparisons with the baseline methods are shown in Tab. 3, demonstrating superior performance in these diverse scenarios and strong generalization capabilities of our proposed method. The sparse-view setting makes the pairwise registration more challenging, thus most of the two-stage multi-view registration baselines fail to achieve good performance. Note that the two learnable multi-view registration baselines [gojcic2020cvpr, wang2023cvpr-sghr] do not work properly with the outdoor LiDAR datasets.
Table 3: Quantitative comparison of multi-view point cloud registration. For each dataset, we report the mean rotation error (RE) in degrees, mean translation error (TE) in meters, and Chamfer distance (CD) in meters. Best results are shown in bold.
3DMatch ScanNet Waymo Scan KAIST Map
Method RE $\downarrow$ TE $\downarrow$ CD $\downarrow$ RE $\downarrow$ TE $\downarrow$ CD $\downarrow$ RE $\downarrow$ TE $\downarrow$ CD $\downarrow$ RE $\downarrow$ TE $\downarrow$ CD $\downarrow$
LMVR [gojcic2020cvpr] 15.46 0.44 0.16 46.10 0.87 0.50 - - - - - -
SGHR [wang2023cvpr-sghr] 50.28 0.78 0.53 23.59 0.64 0.34 - - - - - -
FGR [zhou2016eccv] + PGO 52.81 0.71 0.49 68.80 1.43 0.76 54.60 25.81 3.55 90.07 105.94 41.07
BUFFER-X [seo2025iccv] + PGO 48.16 0.74 0.48 47.63 1.31 0.52 38.37 24.32 2.90 99.47 148.88 39.17
Ours 7.30 0.22 0.10 13.66 0.35 0.12 1.10 1.56 0.23 5.52 9.36 2.97
Furthermore, we showcase the zero-shot generalization ability of our model on the WHU-TLS dataset [dong2020jprs-whutls] in Fig. 4 and on the ETH-TLS dataset [theiler2015jprs] in Fig. 1. Note that our model was never trained on any terrestrial laser scanner (TLS) point clouds, but learned strong geometric priors from the rich outdoor LiDAR data.
Figure 4: Multi-view registration results of our model on WHU-TLS dataset [dong2020jprs-whutls]. Note that our model is never trained on a terrestrial laser scanner point cloud dataset and generalizes zero-shot to this dataset. Left: heritage building scene; Right: park scene. Different colors represent different point clouds. CD refers to Chamfer distance between the registered point cloud and the ground truth point cloud.
Downstream robotics applications. We also showcase in Fig. 5, the potential of our model for downstream robotics applications, including relocalization, multi-robot collaborative mapping, and multi-session map merging. The mix of training data for scan-to-scan, scan-to-map and map-to-map registration allows our model to handle various types of point cloud data collected by robotic platforms.
Figure 5: Our model has the potential to support multiple robotics downstream tasks, including (a) Relocalization: we show the pairwise registration result of two submap point clouds with low overlap in a challenging forest scene from WildPlace [knights2023icra-wildplaces] dataset; (b) Multi-robot collaborative mapping: we show the multi-view registration result of six LiDAR scans in a campus scene from Oxford Spires [tao2025ijrr-oxfordspires] dataset; (c) Multi-session map merging: we show the multi-view registration result of three map-level point clouds of the same place collected from different sessions from KAIST [jung2024ijrr] dataset. Different colors represent different point clouds.

### v2

@@ 3.4 Canonicalized Registration Pipeline
(i) Keypoint selection with local descriptors.  Directly generating millions of points is computationally inefficient and unstable. Instead, we construct a compact, geometry-aware representation by sampling keypoints in each view and attaching local descriptors. For each input point cloud $\mathbf{P}_{i}$, we first perform voxel downsampling with voxel size $v_{d}$, obtaining a reduced cloud $\mathbf{P}_{i}^{v}$. We then apply the farthest point sampling (FPS) to $\mathbf{P}_{i}^{v}$ to select $K_{i}$ keypoints, forming $\mathbf{Q}_{i}\in\mathbb{R}^{3\times K_{i}}$, with the total keypoint count $K:=\sum_{i}K_{i}$ across all views. To obtain uniform coverage over the scene, we choose $K_{i}$ in proportion to the metric scale of $\mathbf{P}_{i}$.
@@
To encode local geometry, for each sampled point in $\mathbf{Q}_{i}$ we define a local patch by a ball query of radius $r_{s}=20v_{d}$ in the reduced point cloud $\mathbf{P}_{i}^{v}$ and extract a local descriptor from the normalized points within this patch. We use the lightweight, rotation-invariant MiniSpinNet [ao2023cvpr-buffer, seo2025iccv] pretrained on 3DMatch [zeng2017cvpr], yielding $\mathbf{F}_{i}\in\mathbb{R}^{32\times K_{i}}$. Concatenating all views, we obtain sampled points $\mathcal{Q}=\{\mathbf{Q}_{i}\in\mathbb{R}^{3\times K_{i}}\}_{i=1}^{N}$ and their local features $\mathcal{F}=\{\mathbf{F}_{i}\in\mathbb{R}^{D\times K_{i}}\}_{i=1}^{N}$, which serve as a compressed yet informative representation of the dense input clouds $\mathcal{P}$.
@@
(ii) Canonicalization of inputs and targets.  To make training invariant to global pose and metric scale, we canonicalize both the conditioning representation and the flow target in a shared frame. For each view $i$, we first translate the sampled points $\mathbf{Q}_{i}$ so that its center of mass is at the origin. We then compute a global scale factor $s$ as the longest edge length of the bounding box of the view with the most points, and scale all centered point sets by $1/s$, so that the entire scene fits in a unit cube. Finally, we apply a random 3D rotation to each centered, scaled cloud. This yields normalized unposed keypoints $\bar{\mathcal{Q}}=\{\bar{\mathbf{Q}}_{i}\in\mathbb{R}^{3\times K_{i}}\}_{i=1}^{N}$ and corresponding similarity transforms $\bar{\mathbf{T}}_{i}\in\mathrm{SIM}(3)$ that map the original $\mathbf{Q}_{i}$ to $\bar{\mathbf{Q}}_{i}$ in the canonical frame.
@@
The training target is defined in the same canonical frame. We first transform the keypoints by the ground-truth poses $\mathcal{\hat{T}}$ to obtain registered keypoints $\mathcal{Q}^{r}=\{\mathbf{Q}_{i}^{r}\}_{i=1}^{N}$ and merge them into $\mathbf{Q}^{r}=\bigcup_{i=1}^{N}\mathbf{Q}_{i}^{r}$. We then (i) recenter $\mathbf{Q}^{r}$ at the origin, (ii) apply the same random rotation used for the view with the most points to fix a reference orientation, and (iii) scale by the global factor $s$, resulting in the normalized registered point cloud $\bar{\mathbf{Q}}^{r}$, which is set as the target $\mathbf{X}(0)$.
@@
(iii) Conditional flow model.  We adopt the Diffusion Transformer [peebles2023iccv] for $\mathbf{V}_{\theta}$, and, following VGGT [wang2025cvpr-vggt], employ a transformer with alternating attention blocks (Fig. 2). Specifically, we alternate per-view self-attention within each point cloud, which consolidates view-specific structure, with global attention over all point tokens to fuse information across views. Our model comprises $L=10$ alternating attention blocks with hidden dimension $d=512$ and $h=8$ attention heads, totaling 73 million parameters.
@@
The model’s condition $\mathbf{C}=f_{\text{emb}}(\bar{\mathcal{Q}},\mathcal{F})$ is obtained via a linear feature embedder $f_{\text{emb}}$ with the concatenated local descriptors $\mathcal{F}$ and the positional encodings of the normalized coordinates $\bar{\mathcal{Q}}$. The flow network $\mathbf{V}_{\theta}$ takes $\mathbf{X}(t)$ and $\mathbf{C}$ as input and is trained with the conditional flow matching loss [lipman2023iclr]. Importantly, we do not condition the model on view indices, making the architecture view-count-agnostic and allowing it to generalize to larger numbers of views at test time than seen during training.
@@
(iv) Lifting to dense registered point clouds.  At inference, the model generates a canonical registered keypoint cloud $\hat{\mathbf{X}}(0)$. Using the rigidity-forcing procedure in Sec. 3.3, we recover per-view rigid transformations $\hat{\mathbf{T}}_{i}$ that align $\bar{\mathbf{Q}}_{i}$ to the corresponding subsets of $\hat{\mathbf{X}}(0)$. The overall transformation from the original input frame of view $i$ to the final registered frame is $\mathbf{T}_{i}=\hat{\mathbf{T}}_{i}\bar{\mathbf{T}}_{i}$, which we apply to all points in the dense cloud $\mathbf{P}_{i}$. Finally, we undo the global scaling by $s$ to obtain the registered point clouds at the metric scale.
@@ 4 Experimental Evaluation
We present our experiments on pairwise and multi-view point cloud registration to show the capabilities of our proposed method named RAP.
@@ 4.1 Experimental Setup
Implementation details.  We train our model for three days with about 120k iterations using 32 NVIDIA A100 GPUs with 80 GB VRAM each. We use Muon [liu2025arxiv-muon] as the optimizer with an initial learning rate of $2\cdot 10^{-3}$ for matrix-like parameters and $2\cdot 10^{-4}$ for vector-like parameters.
Unlike VGGT [wang2025cvpr-vggt] or RPF [sun2025neurips], our training samples have varying numbers of views $N$ and feature point (token) counts $K$. To efficiently train under this irregular data setup, we devise a dynamic batching strategy that allocates samples to each GPU with a suitable batch size. We set the maximum token count per batch on one GPU to 110,000.
@@
For our model, we curate over 100k multi-view registration instances from 17 diverse datasets spanning outdoor, indoor, and object-centric settings. We train on 12 outdoor LiDAR datasets KITTI [geiger2012cvpr], KITTI360 [liao2022pami], Apollo [huang2018cvprws], nuScenes [caesar2020cvpr], MulRAN [kim2020icra], Boreas [burnett2023ijrr-boreas], Oxford Spires [tao2025ijrr-oxfordspires], VBR [brizi2024icra-vbr], UrbanNav [hsu2023navi-urbannav], WildPlace [knights2023icra-wildplaces], HeLiPR [jung2024ijrr], and KITTI-Carla [deschaud2021arxiv-kitticarla]. We also use 4 indoor depth camera datasets 3DMatch [zeng2017cvpr], NSS [sun2025jprs], ScanNet [dai2017cvpr], and ScanNet++ [yeshwanth2023iccv-scannetpp], and the object-centric dataset ModelNet [wu2015cvpr]. Each sample consists of $N$ views of point clouds, where $2\leq N\leq 16$. We split the samples into training and validation sets with an approximate ratio of 9:1 while we also exclude sequences used for testing in common registration benchmarks from the training set. The datasets span diverse scenes across multiple continents, captured by LiDAR and depth cameras with varying resolutions and fields of view. We curate both single-frame and sequence-accumulated submap samples. We require all point clouds in a sample to form a connected overlap graph, but deliberately include hard samples with very low pairwise overlap to improve the model’s robustness to low-overlap scenarios.
4.2 Pairwise Registration Evaluation
We first evaluate our model on pairwise registration across six widely used benchmarks: ModelNet [wu2015cvpr] for object-centric scenes; 3DMatch [zeng2017cvpr], 3DLoMatch [huang2021cvpr-predator], and NSS [sun2025jprs] for indoor depth-camera scenarios; and ETH [pomerleau2012ijrr] and KITTI [geiger2012cvpr] for outdoor LiDAR scenarios. Following prior work, we evaluate performance using registration success rate (%), computed with thresholds on correspondence RMSE for 3DMatch and 3DLoMatch, and on translation and rotation error for the remaining datasets. We adopt the success thresholds used in previous studies [zeng2017cvpr, sun2025jprs, seo2025iccv]; details are provided in the supplementary material. We compare against three conventional baselines based on hand-crafted features (FPFH+FGR [zhou2016eccv], FPFH+TEASER [yang2020tro], KISS-Matcher [lim2025icra-kissmatcher]) and six learning-based methods (FCGF [choy2019iccv], Predator [huang2021cvpr-predator], GeoTransformer [qin2022cvpr-geotransformer], BUFFER [ao2023cvpr-buffer], PARENet [yao2024iccv-parenet], BUFFER-X [seo2025iccv]).
Tab.˜1 shows that our model ranks first on half of the benchmarks and achieves the best average performance across all six benchmarks compared to state-of-the-art methods. We note that our model is trained on more diverse data, while some learning-based baselines are trained solely on 3DMatch or KITTI. We provide additional results on pairwise registration under low overlap in the supplementary material.
Table 1: Pairwise registration success rate (%) on six commonly used pairwise registration benchmarks for object-centric, indoor, and outdoor scenarios. The best result is in bold, and the second best is underscored.
Method ModelNet 3DMatch 3DLoMatch NSS ETH KITTI Avg.
FPFH+FGR [zhou2016eccv] 84.04 62.53 15.42 30.82 91.87 98.74 63.90
FPFH+TEASER [yang2020tro] 86.10 52.00 13.25 25.78 93.69 98.92 61.62
KISS-Matcher [lim2025icra-kissmatcher] 87.12 67.22 20.44 53.69 96.77 100.0 70.87
FCGF [choy2019iccv] 16.51 88.18 40.09 42.86 55.53 98.92 58.68
Predator [huang2021cvpr-predator] 84.36 90.60 62.40 92.99 54.42 99.82 80.77
GeoTransformer [qin2022cvpr-geotransformer] 86.26 92.00 75.00 55.59 71.53 99.82 80.03
BUFFER [ao2023cvpr-buffer] 92.42 92.90 71.80 72.44 99.30 99.64 88.12
PARENet [yao2024iccv-parenet] 66.14 95.00 80.50 45.07 69.42 99.82 75.99
BUFFER-X [seo2025iccv] 99.84 95.58 74.18 85.60 99.72 99.82 92.46
RAP (Ours) 99.13 95.90 78.78 96.27 99.44 100.0 94.92
4.3 Multi-view Registration Evaluation
Figure 3: Overview of our proposed cross-domain multi-view registration benchmark, showing key dataset statistics (top-left), covered application domains (top-right), and representative samples from each scenario category (bottom). Different colors of the point cloud denote different views.
Cross-domain multi-view registration benchmark.  We identify a gap in the evaluation of point cloud registration methods on challenging real-world data. While pairwise registration success rates on established benchmarks such as ModelNet, ETH, and KITTI are nearly saturated (see Tab.˜1), methods that perform well on these benchmarks often fail to generalize to practical applications. Furthermore, while some prior works use indoor depth-camera datasets [zeng2017cvpr, dai2017cvpr, sun2025jprs, choi2015cvpr] for multi-view registration evaluation, no cross-domain benchmark exists covering diverse scales, scenarios, and sensor modalities.
To address this gap, we propose a challenging cross-domain multi-view registration benchmark (Fig.˜3). It aggregates point cloud datasets from diverse sources, most of which were not designed for registration evaluation, and spans five scene categories: object, indoor scan, outdoor scan, terrestrial laser scanning (TLS), and map. Crucially, none of these datasets overlap with our training mixture, making evaluation zero-shot. The benchmark covers practical applications including robot grasping, indoor relocalization, multi-session map merging, collaborative SLAM, TLS surveying, and remote sensing change detection. Dataset sources and curation details are provided in the supplementary material.
Evaluation metrics and baselines.  We adopt the edge success rate and graph success rate as our main evaluation metrics. The edge success rate is the mean over all valid edges (i.e., pairs with non-zero overlap) in a sample, while the graph success rate requires every valid edge in the multi-view registration graph to be successful. To handle diverse scene scales, we normalize translation error by the longest axis of the ground-truth bounding box; a registration succeeds if the normalized translation error is below 2.5% and rotation error below 15°, a threshold reliably correctable by ICP [besl1992pami]. Stricter threshold results (0.5%, 3°) are in the supplementary.
We select the top-3 performing conventional and learning-based pairwise registration methods from Tab.˜1 as baselines, each paired with a robust and efficient implementation of pose graph optimization [choi2015cvpr] to form a complete two-stage multi-view registration pipeline. For a fair comparison, we retrain BUFFER-X, the best-performing baseline in the pairwise evaluation, on the same extended training data as RAP, denoting this variant BUFFER-X. We also include two dedicated multi-view methods: SGHR [wang2023cvpr-sghr], which uses learned overlap scores and history-reweighting iterative optimization, and RPF [sun2025neurips], a flow-matching method targeting object-level shape assembly.
Experiment results.  As shown in Tab.˜2, RAP achieves the best performance across all five scenario categories by a large margin, improving the overall edge success rate from 57.7% to 85.9% and graph success rate from 36.5% to 77.3% compared to the strongest pairwise baseline BUFFER-X. Training BUFFER-X with more diverse data yields a modest gain over the original model, yet still falls significantly short of RAP, indicating that data scaling alone is insufficient without a more capable model architecture and training target. Other learning-based methods (Predator [huang2021cvpr-predator], BUFFER [ao2023cvpr-buffer], SGHR [wang2023cvpr-sghr], RPF [sun2025neurips]) fail to generalize across domains; RPF in particular struggles as it operates on raw points without local descriptors, which does not scale to large scenes. KISS-Matcher [lim2025icra-kissmatcher] is relatively competitive (46.2%/29.2%) thanks to its scalable hand-crafted features and robust transformation estimator, but is still far behind RAP. RAP w/o rigidity forcing (RF) already surpasses all baselines, with test-time rigidity forcing integration providing further consistent improvement.
Notably, RAP achieves the shortest runtime at 8.9 s per sample, outpacing all baselines including KISS-Matcher+PGO (18.4 s), BUFFER-X+PGO (36.0 s), and SGHR (49.7 s). This efficiency stems from RAP’s single-stage design: unlike two-stage pipelines that first perform pairwise registration on all or a selected subset of likely-overlapping pairs (as in SGHR) and then run pose graph optimization, RAP directly generates the registered point cloud in one forward pass without any pairwise stage. We further analyze how runtime scales with view count and token count in Fig.˜5.
Table 2: Comparison of the zero-shot testing performance for multi-view registration on the cross-domain multi-view registration benchmark. We report the registration success rate (%) calculated for edges and graphs (with a threshold of 2.5% for normalized translation error and 15 for rotation error) in different scenarios as well as the average runtime. The best result is in bold, and the second best is underscored.
Method Object Indoor Outdoor TLS Map All Runtime (s)
edge graph edge graph edge graph edge graph edge graph edge graph
FPFH+FGR [zhou2016eccv] +PGO 14.4 6.8 2.8 1.3 4.3 1.1 5.6 2.0 9.0 7.7 5.5 2.2 19.7
FPFH+TEASER [yang2020tro] 18.3 13.5 8.7 5.3 13.9 9.6 17.3 12.4 30.2 19.4 16.7 11.4 22.2
KISS-Matcher [lim2025icra-kissmatcher] 38.1 29.7 30.5 14.0 33.4 19.4 63.9 42.5 81.8 70.3 46.2 29.2 18.4
Predator [huang2021cvpr-predator] 2.8 1.9 9.0 5.9 7.2 5.1 4.3 3.6 8.2 5.7 6.8 4.6 51.8
BUFFER [ao2023cvpr-buffer] 2.3 1.1 25.5 12.0 16.7 9.8 28.7 16.9 3.9 1.5 22.5 11.8 250.3
BUFFER-X [seo2025iccv] 67.6 56.8 43.4 22.3 44.6 31.0 63.9 42.1 26.9 17.2 50.4 32.4 36.0
BUFFER-X* [seo2025iccv] 70.3 59.2 47.3 25.1 54.0 37.1 69.2 46.8 33.6 21.5 57.7 36.5 36.0
SGHR [wang2023cvpr-sghr] 12.6 8.1 28.9 22.0 19.8 11.0 24.6 17.6 15.1 10.9 24.3 17.2 49.7
RPF [sun2025neurips] 2.5 1.7 0.8 0.5 0.7 0.3 0.9 0.3 1.3 0.9 1.0 0.6 11.5
RAP (Ours) 86.3 85.1 75.1 59.6 86.6 80.6 92.2 84.6 94.0 91.6 85.9 77.3 8.9
RAP (Ours) w/o RF 83.1 82.4 72.3 54.7 84.7 76.6 90.1 80.0 89.7 86.8 83.6 72.9 8.9
Figure 4: Multi-view registration performance analysis on the cross-domain multi-view registration benchmark. Left: Edge success rate (mean over pairs) vs. edge overlap ratio. Right: Graph success rate (all pairs must succeed) vs. number of views. RAP outperforms BUFFER-X* and KISS-Matcher across overlap and view counts.
Fig.˜4 provides an in-depth analysis of RAP’s performance as a function of overlap ratio and view count, comparing against the two best-performing baselines. As shown on the left, RAP maintains a consistently higher edge success rate than BUFFER-X+PGO and KISS-Matcher+PGO across all overlap ratios. The performance gap is modest at high overlap but widens substantially as overlap decreases. At an overlap ratio of 20%, RAP still achieves above 80% success rate, while the baselines drop below 40%, demonstrating the superior robustness of our method under low-overlap conditions. The right plot shows that the graph success rate of RAP remains stable and even improves with more views, while the baselines deteriorate more steeply. This highlights a key weakness of two-stage pipelines: erroneous pairwise registrations propagate as outliers into the pose graph, which pose graph optimization cannot reliably resolve. In contrast, our method does not adjust the per-view transformations explicitly. Instead, it performs an implicit global adjustment at the point level, transporting all keypoints across all views simultaneously toward a globally consistent configuration. Additional views therefore provide complementary geometric evidence rather than additional sources of pairwise error.
We provide additional multi-view registration results on 3DMatch [zeng2017cvpr] and ScanNet [dai2017cvpr] in the supplementary material.